A transverse wave of amplitude 0.50 mm and frequency 100 Hz is produced on a wire stretched to a tension of 100 N. If the wave speed is 100 m s-1, what average power is the source transmitting to the wire?

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. r = 0.5 mm = 0.5 * 10–3 mt f = 100 Hz, T = 100 N v = 100 m/s v = T /m ⇒ v2 = (T/m) ⇒ m = (T/v2) = 0.01 kg/m Pave = 2π2 mvr2f2 = 2(3.14)2(0.01) * 100 * (0.5 * 10–3)2 * (100)2 ⇒ 49 * 10–3 watt = 49 mW